OpenStudy (anonymous):

Simplify the expression (-5i)(6i)

OpenStudy (primeralph):

-5*6*i*i = -30*-1 = 30.

OpenStudy (jdoe0001):

recall that \(\bf i^2\implies \sqrt{-1}\cdot \sqrt{-1}\implies (\sqrt{-1})^2\implies \sqrt{(-1)^2}\implies {\color{blue}{ -1}}\)